# MVGT: A Multi-view Graph Transformer Based on Spatial Relations for EEG Emotion Recognition

### v5

@@ 1 Introduction
Affective computing is commonly employed for the analysis of emotional states through Human-Computer Interaction (HCI) systems, which collect multimodal data from subjects, including voice signals, self-reports, body gestures and physiological signals. Among these modalities, physiological signals offer distinct advantages as they are directly derived from the subjects’ mental states, making it difficult for them to disguise or conceal their emotions. Scalp electroencephalography (EEG), a physiological signal usually used to analyze the cognitive functions of the human brain, is collected using noninvasive electrodes on the scalp [alarcao2017emotions]. Nowadays, due to its high temporal resolution, portability, and affordability, this method is widely employed to study brain changes in response to emotional stimuli [niemic2004studies].
@@
EEG signals contain rich emotional information in the temporal, frequency, and spatial domains. How to reasonably and effectively utilize information from each domain is a major challenge. Therefore, recent studies have attempted to combine multi-domain information to capture complex EEG features [MD-AGCN, EmoGT, li2022multi]. Although these works have achieved good performance, their limited use of domain-specific information has constrained model performance.
@@
The frequency domain features of EEG commonly include power spectral density (PSD), differential entropy (DE), differential asymmetry (DASM), and others. Among these, DE has been proven to be the simplest and most effective in EEG emotion recognition tasks [duan2013differential]. DE has the ability to distinguish EEG patterns between high-frequency and low-frequency energy [zheng2015investigating].
@@
Given the high temporal resolution of EEG signals, many studies use temporal models to extract sequential information. Recurrent neural networks (RNNs) have significantly improved emotion recognition performance by modeling temporal dependencies [R2G-STNN, BiDANN]. However, their serial structure poses challenges for parallelization. Convolutional neural networks (CNNs) have also been applied to extract temporal features, but their performance is restricted by limited receptive fields [lggNet, MD-AGCN]. To overcome these difficulties, [EmoGT] employs the attention mechanism to capture relationships between time points; however, this method may suffer performance degradation due to time-unaligned events at a single moment.
@@
The spatial domain of EEG is equally crucial for understanding emotional states, as emotions often involve distributed neural circuits rather than isolated regions [mauss2010measures]. Asymmetric activity between the left and right hemispheres can reflect changes in valence and arousal [schmidt2001frontal]. Message-passing-based graph neural network (MPGNN) methods can effectively capture structural relationships between EEG channels [rgnn, DGCNN, lggNet, MD-AGCN]. However, using MPGNN may pose potential risks of over-smoothing and over-squashing, which could hinder its ability to learn complex representations [ller2024attending]. Additionally, these methods often underutilize the geometric and anatomical structural relationships inherent in EEG spatial information.
@@
To address these limitations and fully exploit features in the temporal, frequency, and spatial domains, we propose a multi-view graph transformer (MVGT) based on spatial relations for EEG emotion recognition. Graph transformers offer robust and flexible spatial perception while mitigating issues like over-smoothing and over-squashing [ller2024attending]. For the frequency domain, we use the embedding of DE as the feature representation. For the temporal domain, inspired by [itransformer], we expand the temporal receptive field by embedding entire continuous time segments as tokens and employ feed-forward neural networks to capture temporal patterns. In the spatial domain, we incorporate several simple yet effective spatial encodings, brain region encoding, centrality encoding and geometric structure encoding, into the model to help it capture effective spatial structures. These encodings allow MVGT to adaptively model correlations between EEG channels, delivering improved performance and deeper insights into emotional states.
@@ 2.2 Temporal Embedding
Temporal information is critical in time-series processing. [EmoGT] treated the features of different channels at the same time points as tokens and employed the attention mechanism to extract temporal correlations between them. However, due to anisotropic volume conduction characteristics [craik2019deep] in human brain tissues, there may be different physical meanings between channels at the same time step, resulting time-unaligned events that subsequently degrade performance. Other approaches [MD-AGCN, lggNet] utilized CNNs to extract temporal information along the time axis from continuous EEG segments, with the receptive field limited by the size of the convolution kernel. Inspired by [itransformer], we broaden the receptive field by embedding entire time segments as tokens rather than single time points. EEG signals are segmented using overlapping sliding windows of size $T$, and these segments are treated as tokens. After processing with sliding windows, we obtain $\tilde{\bm{X}}=\left(\tilde{\bm{X}}_{1},\tilde{\bm{X}}_{2},\cdots,\tilde{\bm{X}}_{S}\right)\in\mathbb{R}^{S\times n\times Tf}$, where $S$ denotes the number of continuous EEG segments, $n$ is the number of channels, and $f$ denotes the dimension of DE features. Refer to the Appendix 5.2 for more temporal embedding details.
According to the universal approximation theorem [hornik1991approximation], the feed-forward neural network (FFN), as the basic module of the Transformer [Transformer] encoder, can learn the intrinsic properties to describe a time series. By processing continuous time segments as input, the FFN extracts temporal features independently for each channel, providing more advantageous predictive representations compared to self-attention [itransformer].
@@ 2.3.1 Brain Region Encoding
Neuroscience research indicates that specific brain regions activate together during high-level cognitive processes [kober2008functional]. In EEG emotion recognition, incorporating relevant neuroscience findings can typically enhance recognition accuracy. For example, studies like [rgnn] and [BiHDM] utilized the asymmetry of neural activity between the left and right hemispheres to enhance recognition performance. [lggNet] divided EEG channels into different regions and combined intra-region convolution with inter-region convolution to capture spatial relationships. Building on [lggNet], we propose four learnable BRE schemes based on neuroscience research, ranging from coarse-grained to fine-grained encodings, namely: LOBE, FRONTAL, GENERAL, and HEMISPHERE (see Appendix 5.4 for design ideas). Coarse-grained region division can aggregate more intra-region information, while fine-grained division can capture more inter-region information. The four schemes mentioned above are shown in Fig. 2.
@@ 2.3.2 Structure Encoding
EEG channels are distributed in the 3D space, and the functional connectivity between them remains imprecisely defined. Therefore, inspired by [graphormer3D], we model the relationships between EEG channels as a fully connected directed graph to avoid making specific assumptions about functional connectivity. Firstly, let $\phi(i,j)$ represent the Euclidean distance between node $i$ and node $j$, and encode $\phi(i,j)$ using a set of Gaussian basis functions [graphormer3D, spinconv]. Let $\bm{b}_{k}\in\mathbb{R}^{n\times n}$ denote one of the Gaussian basis functions. The element $(i,j)$ of this function can be expressed as:
@@
where $\mu_{k}$ (mean), $\sigma_{k}$ (variance), $\alpha_{ij}$, $\beta_{ij}$ are learnable scalars, and $i$ and $j$ denote the index of the source and target node, respectively. By utilizing multiple Gaussian basis functions, we project the distance $\phi$ into various distributions. A distance value close to the mean $\mu_{k}$ of the $k$-th basis function receives higher attention, while the variance $\sigma_{k}$ controls the attention’s concentration. This configuration provides flexibility and expressiveness to the encoding. The outputs of these functions are concatenated, forming the structure encoding:
@@
Unlike previous symmetric adjacency matrices [EmoGT, rgnn], our directed spatial encoding matrix allows the model to learn distinct correlations for $\left(i,j\right)$ and $\left(j,i\right)$. Letting $l$ denote the model depth, $s$ denote the sample index and $m$ denote the index of multi-head attention, the multi-head attention module (MHA) can be represented as:
@@ 2.4 Implementation Details of MVGT
In this section, we describe the overall architecture of the model, as illustrated in Fig. 1. For better optimization, we first apply GraphNorm [graphnorm] to normalize the input features to a range between 0 and 1. The encodings could be characterized as below:
@@
We employ a Pre-LN structure, applying layer normalization (LN) before the MHA and the FFN. This design choice is supported by prior research [xiong2020on], which demonstrates that Pre-LN structures produce more stable gradients, facilitating faster and more reliable convergence. To further prevent overfitting, dropout is employed during training. The process is expressed mathematically as:
@@
Inspired by [graphormer3D, AlphaFold], we apply iterative refinement by recursively feeding the outputs of the model back into the same modules, denoted as “recycling” in Fig. 1. The iterative approach refines the model’s ability to discriminate encoded information and understand emotional patterns, thereby helping the model capture more effective details.
@@ 3.1 Datasets and Settings
For our experiments, we select the SEED [zheng2015investigating] and SEED-IV [seediv] datasets to evaluate the effectiveness of our model. These datasets consist of EEG signals recorded from subjects watching emotion-eliciting videos. Following the settings of previous studies [zheng2015investigating, BiHDM, DGCNN, rgnn, MD-AGCN, EmoGT, seediv, li2022multi], we use pre-computed differential entropy (DE) features for the recognition task. Time segments are extracted using an overlapping sliding window of size $T$, consistent with [MD-AGCN], with $T$ set to 5. For the SEED dataset, we use the first 9 trials of each subject as the training set and the last 6 trials as the test set, as done in previous research. The DE features are computed using five frequency bands extracted from 1s nonoverlapping windows. The model performance is evaluated based on the average accuracy and standard deviation across all subjects over two sessions of EEG data. Similarly, for the SEED-IV dataset, we allocate the first 16 trials to training and the last 8 trials to testing. The DE features for SEED-IV are calculated using 4s windows. The performance of our model is assessed using data from all three sessions. Refer to Appendix 5.1 for DE feature extraction.
During experiments, the hidden dimension is set to 64 and the number of Gaussian basis functions is 32. The number of MHA layers is 4 and the number of attention heads is 2. We employ iterative refinement by recursively feeding the output back into the model for three times. We set the batch size to 32 and the learning rate within the range of 3e-5 to 3e-3. Cross-entropy is used as the loss function, and AdamW [adamw] is employed as the optimizer with a weight decay rate of 0.1.
@@
DGCNN [DGCNN] 90.40/08.49 69.88/16.29
BiDANN [BiDANN] 92.38/07.04 70.29/12.63
BiHDM [BiHDM] 93.12/06.06 74.35/14.09
R2G-STNN [R2G-STNN] 93.34/05.96 -
RGNN [rgnn] 94.24/05.95 79.37/10.54
MD-AGCN [MD-AGCN] 94.81/04.52 87.63/05.77
EmoGT [EmoGT] 95.02/05.99 91.20/09.60
MV-SSTMA [li2022multi] 95.32/3.05 92.82/5.03
@@ 3.4 Visualization of Inter-channel Relations
On the SEED dataset, the parameters suggest that emotional patterns involve activity across multiple brain regions. The parameters from head-1 in MVGT-F suggest that brain activity is primarily concentrated in the temporal and frontal regions on the lateral side of the brain, which aligns with previous research [zheng2015investigating, rgnn, EmoGT]. Notably, the channel connections (TP8, T7), (FT8, T7), and (T8, FT7) reveal a strong correlation between the left and right hemispheres. The connection (FC6, C6) in head-2 shows the strongest activity, further indicating that local information in the central region may play a key role in emotion recognition.
For the SEED-IV dataset, the strongest inter-channel connections occur in the frontal, parietal, and occipital regions, consistent with existing findings [rgnn]. In MVGT-G, the main channel connections are (O1, PO3), (P4, P2), (CB1, PO7), and (PO5, PO7). Additionally, the connection between AF3 and FP1 provides critical insights into emotion processing, in line with [rgnn, EmoGT].
Our model incorporates both intra-regional and inter-regional brain dynamics, highlighting the critical role of distributed network interactions in modeling emotional states, as discussed by [barrett2013large].
@@ 5.3.2 Baseline Models
DGCNN [DGCNN]: A dynamic graph neural network method based on Chebyshev polynomials dynamically learns inter-channel relations in emotion recognition.
BiDANN [BiDANN]: The bi-hemispheres domain adversarial neural network uses recurrent neural networks to capture temporal features and extracts hemisphere information by modeling the left and right hemispheres.
BiHDM [BiHDM]: This model employs a pairwise subnetwork to capture the discrepancy between the left and right hemispheres of the brain.
R2G-STNN [R2G-STNN]: A model that captures spatial-temporal features from local to global scales for emotion classification.
RGNN [rgnn]: A regularized GNN that learns topological relationships between channels.
MD-AGCN [MD-AGCN]: An adaptive GNN that comprehensively considers temporal domain, frequency domain, and brain functional connectivity.
EmoGT [EmoGT]: An elastic graph Transformer network that integrates temporal and spatial information of EEG signals.
MV-SSTMA [li2022multi]: A multi-view masked autoencoder combining CNN and Transformer for emotion recognition.
@@ 5.4 Brain Region Encoding
    Asymmetric EEG activity in the frontal lobe can be utilized for discriminating valence changes [schmidt2001frontal]. The left frontal lobe exhibits a stronger correlation with joy and happiness, while the right frontal lobe is more strongly correlated with fear and sadness. Thus we further divide the frontal lobe region into two symmetrical regions to obtain the FRONTAL scheme.
@@
    According to the symmetry of brain structure [grabner2012oscillatory], we make a finer division of the brain lobe regions, defining the HEMISPHERE scheme.

### v6

@@ 1 Introduction
Affective computing is commonly employed for the analysis of emotional states through Human-Computer Interaction (HCI) systems, which collect multimodal data from subjects, including voice signals, self-reports, body gestures and physiological signals. Among these modalities, physiological signals offer distinct advantages as they are directly derived from the subjects’ mental states, making it difficult for them to disguise or conceal their emotions. Scalp electroencephalography (EEG), a physiological signal usually used to analyze the cognitive functions of the human brain, is collected using noninvasive electrodes on the scalp [1]. Nowadays, due to its high temporal resolution, portability, and affordability, this method is widely employed to study brain changes in response to emotional stimuli [21].
@@
EEG signals contain rich emotional information in the temporal, frequency, and spatial domains. How to reasonably and effectively utilize information from each domain is a major challenge. Therefore, recent studies have attempted to combine multi-domain information to capture complex EEG features [12, 9, 13]. Although these works have achieved good performance, their limited use of domain-specific information has constrained model performance.
@@
The frequency domain features of EEG commonly include power spectral density (PSD), differential entropy (DE), differential asymmetry (DASM), and others. Among these, DE has been proven to be the simplest and most effective in EEG emotion recognition tasks [6]. DE has the ability to distinguish EEG patterns between high-frequency and low-frequency energy [29].
@@
Given the high temporal resolution of EEG signals, many studies use temporal models to extract sequential information. Recurrent neural networks (RNNs) have significantly improved emotion recognition performance by modeling temporal dependencies [15, 16]. However, their serial structure poses challenges for parallelization. Convolutional neural networks (CNNs) have also been applied to extract temporal features, but their performance is restricted by limited receptive fields [5, 12]. To overcome these difficulties, [9] employs the attention mechanism to capture relationships between time points; however, this method may suffer performance degradation due to time-unaligned events at a single moment.
@@
The spatial domain of EEG is equally crucial for understanding emotional states, as emotions often involve distributed neural circuits rather than isolated regions [19]. Asymmetric activity between the left and right hemispheres can reflect changes in valence and arousal [22]. Message-passing-based graph neural network (MPGNN) methods can effectively capture structural relationships between EEG channels [30, 25, 5, 12]. However, using MPGNN may pose potential risks of over-smoothing and over-squashing, which could hinder its ability to learn complex representations [20]. Additionally, these methods often underutilize the geometric and anatomical structural relationships inherent in EEG spatial information.
@@
To address these limitations and fully exploit features in the temporal, frequency, and spatial domains, we propose a multi-view graph transformer (MVGT) based on spatial relations for EEG emotion recognition. Graph transformers offer robust and flexible spatial perception while mitigating issues like over-smoothing and over-squashing [20]. For the frequency domain, we use the embedding of DE as the feature representation. For the temporal domain, inspired by [17], we expand the temporal receptive field by embedding entire continuous time segments as tokens and employ feed-forward neural networks to capture temporal patterns. In the spatial domain, we incorporate several simple yet effective spatial encodings, brain region encoding, centrality encoding and geometric structure encoding, into the model to help it capture effective spatial structures. These encodings allow MVGT to adaptively model correlations between EEG channels, delivering improved performance and deeper insights into emotional states.
@@ 2.2 Temporal Embedding
Temporal information is critical in time-series processing. [9] treated the features of different channels at the same time points as tokens and employed the attention mechanism to extract temporal correlations between them. However, due to anisotropic volume conduction characteristics [4] in human brain tissues, there may be different physical meanings between channels at the same time step, resulting time-unaligned events that subsequently degrade performance. Other approaches [12, 5] utilized CNNs to extract temporal information along the time axis from continuous EEG segments, with the receptive field limited by the size of the convolution kernel. Inspired by [17], we broaden the receptive field by embedding entire time segments as tokens rather than single time points. EEG signals are segmented using overlapping sliding windows of size $T$, and these segments are treated as tokens. After processing with sliding windows, we obtain $\tilde{\bm{X}}=\left(\tilde{\bm{X}}_{1},\tilde{\bm{X}}_{2},\cdots,\tilde{\bm{X}}_{S}\right)\in\mathbb{R}^{S\times n\times Tf}$, where $S$ denotes the number of continuous EEG segments, $n$ is the number of channels, and $f$ denotes the dimension of DE features. Refer to the Appendix 5.2 for more temporal embedding details.
According to the universal approximation theorem [8], the feed-forward neural network (FFN), as the basic module of the Transformer [26] encoder, can learn the intrinsic properties to describe a time series. By processing continuous time segments as input, the FFN extracts temporal features independently for each channel, providing more advantageous predictive representations compared to self-attention [17].
@@ 2.3.1 Brain Region Encoding
Neuroscience research indicates that specific brain regions activate together during high-level cognitive processes [11]. In EEG emotion recognition, incorporating relevant neuroscience findings can typically enhance recognition accuracy. For example, studies like [30] and [14] utilized the asymmetry of neural activity between the left and right hemispheres to enhance recognition performance. [5] divided EEG channels into different regions and combined intra-region convolution with inter-region convolution to capture spatial relationships. Building on [5], we propose four learnable BRE schemes based on neuroscience research, ranging from coarse-grained to fine-grained encodings, namely: LOBE, FRONTAL, GENERAL, and HEMISPHERE (see Appendix 5.4 for design ideas). Coarse-grained region division can aggregate more intra-region information, while fine-grained division can capture more inter-region information. The four schemes mentioned above are shown in Fig. 2.
@@ 2.3.2 Structure Encoding
EEG channels are distributed in the 3D space, and the functional connectivity between them remains imprecisely defined. Therefore, inspired by [23], we model the relationships between EEG channels as a fully connected directed graph to avoid making specific assumptions about functional connectivity. Firstly, let $\phi(i,j)$ represent the Euclidean distance between node $i$ and node $j$, and encode $\phi(i,j)$ using a set of Gaussian basis functions [23, 24]. Let $\bm{b}_{k}\in\mathbb{R}^{n\times n}$ denote one of the Gaussian basis functions. The element $(i,j)$ of this function can be expressed as:
@@
where $\mu_{k}$ (mean), $\sigma_{k}$ (standard deviation), $\alpha_{ij}$, $\beta_{ij}$ are learnable scalars, and $i$ and $j$ denote the index of the source and target node, respectively. By utilizing multiple Gaussian basis functions, we project the distance $\phi$ into various distributions. A distance value close to the mean $\mu_{k}$ of the $k$-th basis function receives higher attention, while the standard deviation $\sigma_{k}$ controls the attention’s concentration. This configuration provides flexibility and expressiveness to the encoding. The outputs of these functions are concatenated, forming the structure encoding:
@@
Unlike previous symmetric adjacency matrices [9, 30], our directed spatial encoding matrix allows the model to learn distinct correlations for $\left(i,j\right)$ and $\left(j,i\right)$. Letting $l$ denote the model depth, $s$ denote the sample index and $m$ denote the index of multi-head attention, the multi-head attention module (MHA) can be represented as:
@@ 2.4 Implementation Details of MVGT
In this section, we describe the overall architecture of the model, as illustrated in Fig. 1. For better optimization, we first apply GraphNorm [3] to normalize the input features to a range between 0 and 1. The encodings could be characterized as below:
@@
We employ a Pre-LN structure, applying layer normalization (LN) before the MHA and the FFN. This design choice is supported by prior research [27], which demonstrates that Pre-LN structures produce more stable gradients, facilitating faster and more reliable convergence. To further prevent overfitting, dropout is employed during training. The process is expressed mathematically as:
@@
Inspired by [23, 10], we apply iterative refinement by recursively feeding the outputs of the model back into the same modules, denoted as “recycling” in Fig. 1. The iterative approach refines the model’s ability to discriminate encoded information and understand emotional patterns, thereby helping the model capture more effective details.
@@ 3.1 Datasets and Settings
For our experiments, we select the SEED [29] and SEED-IV [28] datasets to evaluate the effectiveness of our model. These datasets consist of EEG signals recorded from subjects watching emotion-eliciting videos. Following the settings of previous studies [29, 14, 25, 30, 12, 9, 28, 13], we use pre-computed differential entropy (DE) features for the recognition task. Time segments are extracted using an overlapping sliding window of size $T$, consistent with [12], with $T$ set to 5. For the SEED dataset, we use the first 9 trials of each subject as the training set and the last 6 trials as the test set, as done in previous research. The DE features are computed using five frequency bands extracted from 1s nonoverlapping windows. The model performance is evaluated based on the average accuracy and standard deviation across all subjects over two sessions of EEG data. Similarly, for the SEED-IV dataset, we allocate the first 16 trials to training and the last 8 trials to testing. The DE features for SEED-IV are calculated using 4s windows. The performance of our model is assessed using data from all three sessions. Refer to Appendix 5.1 for DE feature extraction.
During experiments, the hidden dimension is set to 64 and the number of Gaussian basis functions is 32. The number of MHA layers is 4 and the number of attention heads is 2. We employ iterative refinement by recursively feeding the output back into the model for three times. We set the batch size to 32 and the learning rate within the range of 3e-5 to 3e-3. Cross-entropy is used as the loss function, and AdamW [18] is employed as the optimizer with a weight decay rate of 0.1.
@@
DGCNN [25] 90.40/08.49 69.88/16.29
BiDANN [16] 92.38/07.04 70.29/12.63
BiHDM [14] 93.12/06.06 74.35/14.09
R2G-STNN [15] 93.34/05.96 -
RGNN [30] 94.24/05.95 79.37/10.54
MD-AGCN [12] 94.81/04.52 87.63/05.77
EmoGT [9] 95.02/05.99 91.20/09.60
MV-SSTMA [13] 95.32/3.05 92.82/5.03
@@ 3.4 Visualization of Inter-channel Relations
On the SEED dataset, the parameters suggest that emotional patterns involve activity across multiple brain regions. The parameters from head-1 in MVGT-F suggest that brain activity is primarily concentrated in the temporal and frontal regions on the lateral side of the brain, which aligns with previous research [29, 30, 9]. Notably, the channel connections (TP8, T7), (FT8, T7), and (T8, FT7) reveal a strong correlation between the left and right hemispheres. The connection (FC6, C6) in head-2 shows the strongest activity, further indicating that local information in the central region may play a key role in emotion recognition.
For the SEED-IV dataset, the strongest inter-channel connections occur in the frontal, parietal, and occipital regions, consistent with existing findings [30]. In MVGT-G, the main channel connections are (O1, PO3), (P4, P2), (CB1, PO7), and (PO5, PO7). Additionally, the connection between AF3 and FP1 provides critical insights into emotion processing, in line with [30, 9].
Our model incorporates both intra-regional and inter-regional brain dynamics, highlighting the critical role of distributed network interactions in modeling emotional states, as discussed by [2].
@@ 5.3.2 Baseline Models
DGCNN [25]: A dynamic graph neural network method based on Chebyshev polynomials dynamically learns inter-channel relations in emotion recognition.
BiDANN [16]: The bi-hemispheres domain adversarial neural network uses recurrent neural networks to capture temporal features and extracts hemisphere information by modeling the left and right hemispheres.
BiHDM [14]: This model employs a pairwise subnetwork to capture the discrepancy between the left and right hemispheres of the brain.
R2G-STNN [15]: A model that captures spatial-temporal features from local to global scales for emotion classification.
RGNN [30]: A regularized GNN that learns topological relationships between channels.
MD-AGCN [12]: An adaptive GNN that comprehensively considers temporal domain, frequency domain, and brain functional connectivity.
EmoGT [9]: An elastic graph Transformer network that integrates temporal and spatial information of EEG signals.
MV-SSTMA [13]: A multi-view masked autoencoder combining CNN and Transformer for emotion recognition.
@@ 5.4 Brain Region Encoding
    Asymmetric EEG activity in the frontal lobe can be utilized for discriminating valence changes [22]. The left frontal lobe exhibits a stronger correlation with joy and happiness, while the right frontal lobe is more strongly correlated with fear and sadness. Thus we further divide the frontal lobe region into two symmetrical regions to obtain the FRONTAL scheme.
@@
    According to the symmetry of brain structure [7], we make a finer division of the brain lobe regions, defining the HEMISPHERE scheme.